OpenStudy (anonymous):

Area of Circle is 484square cm. compute the length of diameter of the circle.

OpenStudy (anonymous):

The formula for the area of a circle is \[A=(\pi)(r^2)\] We know that A=484 Now, we need to find the radius first, since its the only logical piece of information we need to know and find out. So in order to isolate it, we divide both sides by pi. When that is done, we are left with \[484/\pi=r^2\] So we now need to find the value of r, so we have to square root both sides so the r is not being squared. That results in \[\sqrt{(484/\pi)}=r\] So when you calculate the answer, just multiply the radius by 2 to find out the diameter of the circle. Complete. Simple enough. :)

OpenStudy (anonymous):

yes thats the anwer.